### v1

# Broadly Tunable Quantum Enhanced Raman Microscopy for Advancing Bioimaging

Broadly Tunable Quantum Enhanced Raman Microscopy for Advancing Bioimaging
††journal: opticajournal††articletype: Research ArticleStimulated Raman scattering (SRS) microscopy has emerged as a powerful technique for probing the spatiotemporal dynamics of molecular bonds with exceptional sensitivity, resolution, and speed. However, classically, its performance remains fundamentally constrained by optical shot noise, which imposes a strict limit on detection sensitivity and speed. Here, we demonstrate a quantum-enhanced SRS microscopy platform that circumvents this barrier by harnessing amplitude-squeezed light. Specifically, we generate a Stokes beam with of amplitude squeezing using traveling-wave optical parametric amplification in second-order nonlinear waveguides, and combine it with a tunable coherent pump to access vibrational modes spanning from to . Applied to quantum imaging of metabolites in biological tissue (pork muscle), our quantum-enhanced Raman microscope achieves an average noise suppression of and a enhancement in signal-to-noise ratio (SNR) – to the best of our knowledge, the largest improvement reported to date in quantum-enhanced SRS microscopy of biological samples.
1 Introduction
Quantum-enhanced optical sensing exploits nonclassical states of light to overcome the fundamental shot-noise limit, thereby enabling measurement sensitivities that surpass the capabilities of conventional optical methodologies [1, 2]. Among nonclassical states of light, squeezed states constitute a particularly powerful resource, as they can redistribute quantum fluctuations to suppress noise in the quadrature of interest, thereby enhancing the signal-to-noise ratio (SNR) in precision measurements [3, 4]. This capability allows the detection of extremely weak signals that would otherwise be obscured by classical noise limits. The potential of optical quantum-enhanced sensing has been demonstrated in several landmark experiments, most prominently through the deployment of squeezed light in gravitational-wave observatories [5, 6]. In microscopy and spectroscopy, such quantum-enhanced approaches open new avenues for probing molecular information at lower analyte concentrations, under reduced illumination intensities, or at accelerated acquisition [7, 8, 9].
Raman-based microscopy provides a powerful approach for probing molecular vibrations with high chemical specificity in a label-free and non-destructive manner [10]. However, the intrinsically weak nature of spontaneous Raman scattering limits both its sensitivity and acquisition speed, particularly in live-cell and clinical imaging contexts. To overcome these challenges, stimulated Raman scattering (SRS) microscopy was developed as a technique that resonantly drives molecular vibrations using synchronized pump and Stokes beams whose frequency difference is tuned to a specific vibrational mode of interest. By amplifying Raman signals by several orders of magnitude while preserving a linear dependence on molecular concentration, SRS enables quantitative, high-speed, and background-free chemical imaging [11, 12]. These capabilities have established SRS as a robust technique for spatially and chemically resolved imaging of lipids, proteins, metabolites, and pharmaceutical compounds in living systems, as well as for real-time biomedical applications, including intraoperative diagnostics [13, 14].
Extending the principles of quantum-enhanced sensing to Raman-based microscopy provides a direct means to overcome the shot-noise limit that currently constrains SRS sensitivity. In conventional implementations, detection can be improved by increasing optical power or extending integration time, but for fragile biological tissues, excitation must remain below photodamage thresholds to preserve their native dynamics [15, 16]. By employing squeezed probe states, which intrinsically exhibit noise levels below the standard shot-noise limit, quantum-enhanced Raman microscopy achieves higher sensitivity and faster acquisition, hereby reduced photodamage [17].
One of the first experimental proof-of-concept demonstrations of quantum-enhanced SRS (QE-SRS) employed continuous-wave amplitude-squeezed light, achieving a noise reduction of 3.6 dB at 2950 cm-1 in polymer samples [18]. Subsequently, pulsed amplitude squeezing enabled sub-shot-noise imaging of molecular bonds within live yeast cells, yielding a 1.3 dB noise reduction at 2850 cm-1 [19]. Additionally, a quantum-enhanced balanced detection (QE-BD) scheme extended operation into the high-power regime (>30 mW), demonstrating a 3.11 dB noise reduction compared to classical detection strategies [20]. Beyond sensitivity improvements, squeezed light has also been leveraged to enhance imaging capabilities [21], while integration of bright squeezed probe beams into live-cell SRS microscopy enabled 20% faster imaging with reduced photodamage [22]. All these studies demonstrate remarkable improvements offered by QE-SRS compared to conventional implementations. However, their investigation range has so far been restricted to a narrow bandwidth of Raman vibrational modes, primarily within the functional group region (2800–3100 cm-1). This limitation restricts the extraction of molecular information, as the Raman signatures of lipids, proteins, and nucleic acids exhibit substantial spectral overlap within this vibrational region. By extending access into the fingerprint region (500–1800 cm-1), it becomes possible to resolve these overlapping contributions and unambiguously distinguish between the vibrational signatures of different biomolecular species, thereby providing a far richer and more specific chemical characterization of biological tissues.
In this work, we present a quantum-enhanced SRS microscope that employs a bright amplitude-squeezed Stokes beam as the probe field. The squeezed beam provides more than 5 dB of noise reduction while maintaining an optical power of 3.75 mW, making it compatible with picosecond SRS imaging. Integrated into a tunable SRS platform, it enables broadband operation across both the fingerprint region (1450–1650 cm-1) and the CH-stretch region (2800–3100 cm-1), allowing chemically specific mapping of polystyrene reference samples and biological tissue with enhanced sensitivity. Whereas previous quantum-enhanced SRS demonstrations were confined to the CH-stretch region, the present work achieves, for the first time, quantum-enhanced vibrational imaging spanning both spectral regions, thereby enabling chemically distinguishable imaging in biologically relevant environments.
2 Stimulated Raman microscopy setup
2.1 Full scheme
Fig. 1 shows the block diagram of the main parts of our quantum-enhanced stimulated Raman scattering (QE-SRS) microscope. The primary light source is a picoEmerald laser (APE Berlin), which delivers picosecond pulsed light suitable for accessing a wide range of Raman vibrational modes. It provides a broadly tunable Raman pump beam (690–970 nm) with up to 1000 mW of average optical power in the whole range, generated by an internal optical parametric oscillator (OPO), and a fixed-wavelength Stokes (coherent) beam at 1064 nm. Both beams have a pulse duration of 11 ps (full width at half maximum) and 80 MHz repetition rate. The Raman pump beam is separated from the 1064 nm Stokes beam by a dichroic mirror (DM, Thorlabs DMLP1000). Subsequently, amplitude modulation at is applied to the Raman pump beam using a combination of polarizing beam splitters (PBS) and an electro-optical modulator (EOM, Thorlabs EO-AM-R-20-C1). The 1064 nm Stokes beam is directed to the squeezer to generate amplitude-squeezed light. The squeezer module enables switching between classical and squeezed-light configurations for the Stokes beam. The modulated Raman pump is then recombined with the amplitude-squeezed Stokes beam on another dichroic mirror DM. After recombination, both beams are directed into a custom-built microscope system. The microscope includes two 50 microscope objectives (MO 50, Thorlabs LMH-50X-1064), which focus the Stokes beam to a spot size of approximately 1.6 m. Each objective lens exhibits approximately 98 % transmission at 1064 nm. For precise and reliable sample positioning, we employ a custom-designed 3D raster-scanning stage (based on 8MTF-200-B43-MEN4-LEN1-025, Standa). An integrated white-light source and a wide-field imaging camera provide straightforward visualization and alignment of the sample prior to quantum-enhanced SRS measurements.
Following interaction with the sample, the Raman pump and Stokes beams are collected by the second MO 50. The Stokes beam is then separated from the Raman pump using a combination of a dichroic mirror (DM, Thorlabs DMSP1020B) and a long-pass filter (LPF, Thorlabs FELH1000) and directed onto a custom-built photodetector (InGaAs PD), adapted from the design of Ref. [19]. The detector uses an InGaAs PIN Hamamatsu G10899-003K photodiode with a measured quantum efficiency of 97 %. The detector is resonantly tuned to amplify signals at the 19.3 MHz modulation frequency. The SRS signal is acquired using a digital lock-in amplifier implemented on the Moku:Lab platform (MOKU, Liquid Instruments), with a 12-bit vertical resolution, and detection bandwidth of 2 kHz around the demodulation frequency achieving low-noise signal recovery. In addition, the Moku:Lab is used to generate the 19.3 MHz drive signal for the EOM, which enables precise amplitude modulation of the pump beam for SRS detection. For synchronization between the XYZ scanning stage and data acquisition, we implement a custom-made Python script, achieving a scan rate of 15 data points per second. While sufficient for proof-of-concept measurements, this rate is modest compared to commercial high-speed SRS systems. The limitation arises from the inherent speed constraints of our mechanical scanning stage, which are slower than resonant scanners or galvo mirrors, and from potential bottlenecks in data transfer and storage associated with our Python-based acquisition.
2.2 Preparation of amplitude-squeezed light
To obtain an amplitude-squeezed Stokes beam, we employed a single-pass optical parametric amplifier (OPA) based on a periodically poled lithium niobate (PPLN) waveguide [23, 24]. A simplified schematic of the source is shown in Fig. 1 (the yellow box). The squeezed-vacuum (SV) state was generated by pumping the PPLN waveguide with a 532 nm beam. To realize an amplitude-squeezed state, the SV was displaced by a coherent beam (displacing beam) using an asymmetric beam splitter composed of a half-wave plate (HWP) and a polarizing beam splitter (PBS). For stable displacement, the relative phase between the SV and the displacing beam was actively stabilized using an auxiliary reference beam introduced into the squeezer module, which provided the error signal for a feedback loop driving active phase control. As a result, we achieved a 5.2 0.2 dB reduction of amplitude quadrature noise with a displacement of 3.75 mW using an OPA pump power of 2.5 mW. A detailed description of the optical layout, phase-stabilization scheme, and measurement procedure is provided in Supplementary 1, Section 1.
We validated the noise-reduction performance of the prepared amplitude-squeezed state by an SRS measurement. Using a simplified experimental setup, we recorded the SRS signal from a 1 mm-thick polystyrene layer at 3050 cm-1, corresponding to the vibrational modes of the C–H bonds within aromatic rings and the associated ring structures. This measurement yielded an improvement of approximately 4 dB in the background noise level, representing one of the highest reported measured enhancements for QE-SRS on polystyrene [18]. Further experimental details are provided in Supplementary 1, Section 2.
2.3 Tunability of QE-SRS microscope
In addition to the high detected squeezing level, our system provides broad spectral tunability, allowing access to both the fingerprint and functional group regions. This broad tunability is critical for biochemical fingerprinting because the fingerprint and CH-stretch regions contain distinct vibrational markers associated with different molecular species. As an example of the tunability, Fig. 2 presents SRS imaging of a polystyrene microparticle obtained using a coherent (classical) Stokes beam. Polystyrene was selected as a well-established reference material owing to its strong Raman signals at well defined Raman shifts across a broad spectral range [25].
We investigated three of the most prominent Raman lines of polystyrene: 3050 cm-1, corresponding to the aromatic C–H stretching mode; 1600 cm-1, associated with the skeletal C=C stretching mode of the phenyl ring; and 1003 cm-1, corresponding to the symmetric ring-breathing mode of the phenyl group. To match the targeted Raman lines, the pump wavelength from the picoEmerald laser was tuned from 803.5 nm to 961 nm, covering the spectral range from the C–H stretching mode to the ring-breathing mode. The Raman pump power was set to approximately 34 mW, while the coherent Stokes beam power was maintained at around 3.75 mW to ensure a sufficient signal level. At each scanning point, the signal was averaged over 1024 acquisitions, resulting in a dwell time of around 66 ms per pixel. The SRS images obtained with our home-built microscopy system show good agreement with the corresponding white-light wide-field image, confirming both the spatial resolution and spectral tunability of the system.
3 Quantum-enhanced SRS bioimaging
3.1 QE-SRS on a pork muscle tissue
To demonstrate the applicability of our QE-SRS platform for biological imaging, we investigated a pork muscle tissue sample. The tissue was sectioned into 10 m-thick slices and mounted on a cover glass. The sample was sealed with water using an additional cover glass, forming a cuvette-like structure that preserved hydration during imaging. This configuration also facilitated heat dissipation induced by optical illumination and improved the structural integrity of the prepared sample. Moreover, the gradual transition in refractive index (glass sample water glass) helped suppress optical scattering typically arising from refractive index mismatches in biological tissues.
The Raman band at 2940 cm-1 originates from CH2/CH3 stretching modes present in both proteins and lipids and therefore serves as an indicator of overall protein–lipid content in tissue. The measurement was performed using an average optical power of 32 mW for the Raman pump and 3.75 mW for the Stokes beam focused on the sample. The white-light wide-field image is shown in Fig. 3(a). Spatial mapping of the muscle structure was achieved via raster scanning with a step size of 750 nm, with signal averaging over 1024 acquisitions per pixel. The measured SRS signal (Fig. 3(b)) clearly reproduces the structural features of the muscle tissue. Fig. 3(d),(e) further illustrate the enhancement in SNR achieved by employing amplitude-squeezed light. In the case of SRS gain detection, the SNR is [18]
where is the SRS signal intensity, and are pump and Stokes beam intensities, respectively, and is the variance of the detected Stokes beam intensity that quantifies the noise level. By using amplitude-squeezed light, we are able to surpass the shot-noise level, thereby enhancing the SNR in proportion to the inverse square root of the noise reduction factor. In Fig. 3(d) and (e), we observed an SNR improvement of approximately 43% across the entire investigated area (m2), correspond to a noise reduction of compared to the coherent Stokes beam configuration. The initial 5.2 dB of squeezing was reduced to 3.3 dB at the detector due to 24% optical loss arising from optics and sample transmission.
In addition, we investigated the scaling behavior of the SNR with respect to the Raman pump power. As discussed above, the SNR is expected to scale linearly with pump power. To validate this, we measured the SNR of the SRS signal within the white square region shown in Fig. 3(b). The results are presented in Fig. 3(c). Here, the SNR was averaged over a m2 area, while the pump power was varied from 12 mW to 49 mW with the Stokes beam power fixed at 3.75 mW. The blue and red dashed lines correspond to linear fits of the acquired data for coherent (red trace) and amplitude-squeezed (blue trace) Stokes beams, respectively. The increased slope observed in the squeezed-light case corresponds to an estimated noise reduction of approximately 3.3 dB, yielding a 43% improvement in SNR, which is consistent with the enhancements observed in panels (d) and (e).
3.2 Multifrequency biological imaging
To demonstrate chemical specificity, we performed SRS imaging at four vibrational bands characteristic of proteins and lipids. Explicitly, we investigated several prominent Raman vibrational modes: (general C–H stretching in proteins, lipids, and nucleic acids), widely used as a marker of total protein and lipid content; (symmetric CH2 stretching in lipids); (amide I band, primarily C=O stretching in proteins), a key indicator of protein content and secondary structure; and (CH2/CH3 bending in lipids and proteins), often used as a marker for lipid–protein balance. Together, these four bands give a complementary biochemical fingerprint of muscle tissue. For these measurements, the Raman pump was tuned to , , , and with corresponding powers at the sample of 32 mW, 32 mW, 30 mW, and 38 mW, respectively. The Stokes beam power was maintained at in all cases. However, systematically higher SNR were observed at the CH-stretch bands (2850 and 2940 cm-1) compared to the fingerprint modes. This enhancement is attributed to the larger Raman cross-sections and the higher abundance of CH groups in tissue lipids and proteins, particularly in lipid-rich structures such as cell membranes and adipose tissue. In contrast, Fig. 4 shows regions with SRS signals at 1450 cm-1 and 1650 cm-1, where the CH-stretch bands at 2850 and 2940 cm-1 are absent. This spectral pattern indicates regions with elevated protein content and minimal lipid contribution, consistent with lean myofibrillar or connective tissue in muscle [26, 27]. The results presented in Fig. 4 demonstrate a consistent enhancement in SNR across all examined Raman shifts when utilizing the amplitude-squeezed Stokes beam. A noise reduction of dB was achieved across all measured vibrational modes, corresponding to a improvement in SNR compared to the classical (coherent) Stokes configuration. Fig. 4(f) and (g) show cross-sections of the SNR maps. These cross-sections clearly illustrate the improvements achieved with squeezed light compared to coherent illumination. This consistent enhancement underscores the robustness of quantum-assisted amplification in improving Raman signal quality. The observed gains highlight the potential of QE-SRS to extend the sensitivity of vibrational imaging in biological tissues, enabling more precise and reliable real-time biochemical analysis, even at lower optical powers or in samples with weak Raman signatures.
4 Conclusion
In summary, we have implemented quantum-enhanced broadband SRS microscopy for biological applications. To this end, we have first generated a pulsed squeezed vacuum beam using single-pass OPA in a PPLN waveguide. By combining the squeezed vacuum with a coherent displacement beam under carefully optimized spatial, temporal, and polarization overlap, we obtained a 5.2 dB amplitude-squeezed bright beam, providing a stable and controllable nonclassical light source for the use as the Stokes beam for QE-SRS microscopy. Together with the tunable Raman pump, the squeezed Stokes beam provided quantum-enhanced SRS microscopy. The setup demonstrated reliable operation with tunability across both the fingerprint (1450–1650 cm-1) and functional group (2800–3100 cm-1) regions, covering many of the most relevant molecular vibrational modes in biological systems.
We validated the performance of the quantum-enhanced SRS microscope by conducting measurements on both benchmark polystyrene samples and biological samples (pork muscle tissue). The biological measurements revealed an average noise reduction of 3.6 dB, corresponding to a 51% improvement in SNR compared to classical detection. To the best of our knowledge, this represents the highest reported enhancement in QE-SRS microscopy applied to biological samples. Importantly, the results confirm that quantum squeezing can be practically integrated into a picosecond-pulsed SRS platform and translated from proof-of-concept demonstrations on simple materials to more complex biological tissue environments. Along with strong squeezing, we also report, for the first time, QE-SRS across both regions in the fingerprint and high-wavenumber CH-stretch regions. Our approach provides quantum SNR enhancement in both regions, enabling sensitive and label-free imaging and identification of important biomarkers, such as proteins, lipids, amid and nucleic acids in biological samples.
Looking forward, further improvements in coupling efficiency, waveguide engineering, and pump power stabilization are expected to increase the achievable squeezing levels and long-term stability of the source. Integration with more advanced lock-in detection schemes and fast-scanning microscopes could enable real-time quantum-enhanced imaging of living biological samples, extending the method from proof-of-concept demonstrations towards practical biomedical applications. Moreover, combining squeezed-light excitation with computational imaging and machine learning approaches may open new possibilities for super-resolved Raman microscopy with reduced photo damage and enhanced molecular contrast.
Acknowledgment
This research was funded by the Danish Agency for Institutions and Educational Grants and QuRaman project under QuantERA supported by Innovation Fund Denmark (1116-00003B), the German Federal Ministry of Education and Research under the grant number 13N16359, the Danish National Research Foundation (bigQ, DNRF142) and the Novo Nordisk Foundation (CBQS, NNF24SA0088433).
Disclosures
The authors declare no conflicts of interest.
Data availability
Data underlying the results presented in this paper are not publicly available but may be obtained from the authors upon reasonable request.
Supplemental document
See Supplementary 1 for supporting content.